Suppose that Daniel has a 2.50 L bottle that contains a mixture of O2, N2, and CO2 under a total pressure of 4.70 atm. He knows that the mixture contains 0.23 mol N2 and that the partial pressure of CO2 is 0.250 atm. If the temperature is 273 K, what is the partial pressure of O2?
Answer #1

Pressure due to nitrogen=(nN2*RT)/V

Pressure due to nitrogen=(0.23 mol*0.0821Latm mol-1 K-1*273K)/2.5L

Pressure due to nitrogen=2.062 atm

Pressure due to O2=total pressure-pressure of N2-pressure of CO2

Pressure due to O2=4.70-2.062-0.250 = 2.388 atm